### v2

@@ 4.2 Semantic Segmentation
Experimental Setting. We use MMSegmentation [76] to conduct our experiments on semantic segmentation. All our experiments are carried out on the ADE20K [77] dataset, testing two algorithms: UperNet [78] and FPN [79]. We refrain from using pre-training and instead train the models from scratch on the dataset. Regarding evaluation metrics, we assess the models based on aAcc and mIoU in 10 high-score categories. In the ablation study of the backbone network, we examine various popular variants of ViT and CNN and explore their performance during 0 - 20,000 iterations.
Segment on ADE20K. We conducted experiments on multiple backbone networks on ADE20K
Table 6: Semantic Segmentation of Vanilla ViT and LKCA on ADE20K Dataset
Backbone Iteration Algorithm aAcc(%) mIoU(%) # Params Flops
ViT-B 2000 UPerNet 40.90 22.28 0.144G 0.108T
LKCA-B 2000 UPerNet 45.48 24.65 0.102G 0.115T
ViT-B 4000 UPerNet 47.41 28.08 0.144G 0.108T
LKCA-B 4000 UPerNet 52.67 30.00 0.102G 0.115T
ViT-B 8000 UPerNet 53.43 31.73 0.144G 0.108T
LKCA-B 8000 UPerNet 50.79 33.73 0.102G 0.115T
ViT-B 16000 UPerNet 61.55 39.33 0.144G 0.108T
LKCA-B 16000 UPerNet 65.49 40.60 0.102G 0.115T
@@
Table 6 provides performance metrics for the UPerNet algorithm using different backbones (ViT and LKCA), iteration algorithms, and iteration counts (2000, 4000, 8000, 16000). At each iteration count, LKCA-B consistently outperforms ViT-B in terms of aAcc and mIoU, indicating superior performance of LKCA-B in the UPerNet task. As the iteration count increases, both models show an improvement in performance, but at each iteration count, LKCA-B maintains a leading position. Overall, LKCA-B demonstrates superior performance compared to ViT-B at different iteration counts, with higher pixel accuracy and mean Intersection over Union. Additionally, LKCA-B exhibits advantages in terms of parameter count and computational complexity.
Table 7: Semantic Segmentation of Various ViT Variants and LKCA on ADE20K Dataset
Backbone Iteration Algorithm aAcc(%) mIoU(%) # Params Flops
Vit-S 10000 UPerNet 55.55 32.92 57.66M 67.25G
ResNet-50 10000 UPerNet 58.12 35.35 64.12M 0.238T
Twins_PCPVT-S 10000 UPerNet 52.56 29.64 53.30M 0.234T
Swin-S 10000 UPerNet 57.18 35.30 51.90M 0.231T
LKCA-S 10000 UPerNet 60.16 37.16 50.58M 65.94G
Vit-S 10000 FPN 56.56 35.84 37.23M 76.14G
Twins_PCPVT-S 10000 FPN 56.74 33.09 33.43M 49.74G
LKCA-S 10000 FPN 62.55 39.03 30.18M 77.92G
Vit-S 20000 UPerNet 58.51 36.61 57.66M 67.25G
ResNet-50 20000 UPerNet 62.76 37.84 64.12M 0.238T
Twins_PCPVT-S 20000 UPerNet 58.91 38.00 53.30M 0.234T
Swin-S 20000 UPerNet 58.00 37.57 51.90M 0.231T
LKCA-S 20000 UPerNet 64.37 40.09 50.58M 65.94G
Vit-S 20000 FPN 63.36 41.64 37.23M 76.14G
Twins_PCPVT-S 20000 FPN 58.51 36.61 33.43M 49.74G
LKCA-S 20000 FPN 67.21 41.81 30.18M 77.92G
@@
Table 7 presents a performance comparison among different backbones (ViT-S, ResNet-50, Swin-S, Twins PCPVT-S, LKCA-S) under various iteration counts and across two different task settings (UPerNet and FPN). In the UPerNet task, LKCA-S consistently outperforms other models at different iteration counts, exhibiting higher aAcc and mIoU. It achieves superior performance compared to other models. Similarly, in the FPN task, LKCA-S demonstrates excellent performance across different iteration counts, showcasing higher aAcc and mIoU, making it competitive against other models. LKCA-S consistently exhibits outstanding performance across different tasks and iteration settings, while maintaining relatively lower computational complexity and parameter counts.
5 Conclusion
@@
In this study, we re-examine the relationship between the attention mechanism in visual transformers and the large kernel convolutional networks, proposing a new visual attention called Large Kernel Convolutional Attention (LKCA). It employs a single large kernel convolution to simplify attention operations. In comparison to the previous approach, Large Kernel Attention (LKA), which uses multiple small kernel convolutions, our method achieves the same effects but is more straightforward and simpler. LKCA combines the advantages of convolutional neural networks and visual transformers, featuring a larger receptive field, locality, and parameter sharing. We explained the superiority of LKCA from the perspectives of both convolution and attention, providing equivalent code implementations for each view. Experimental results demonstrate that LKCA, implemented from both convolutional and attention perspectives, exhibits comparable performance. We extensively experimented with LKCA variants of ViT on multiple benchmark datasets, including CIFAR-10, CIFAR-100, SVHN, Tiny-ImageNet, and ADE20K, for classification and segmentation tasks. The experimental outcomes reveal that LKCA demonstrates competitive performance in visual tasks.

### v3

@@ IV-D Semantic Segmentation Results
In this work, we follow prior works to choose image classification as the main task for evaluating the efficacy of model architecutres. Nonetheless, we additionally provide results of semantic segmentation on the ADE20K [81] dataset. Specifically, we use MMSegmentation [82] to conduct our experiments. All our experiments are testing two algorithms: UperNet [83] and FPN [84]. We refrain from using pre-training and instead train the models from scratch on the dataset. Regarding evaluation metrics, we assess the models based on aAcc and mIoU in 10 high-score categories. In the ablation study of the backbone network, we examine various popular variants of ViT and CNN and explore their performance during 0 - 20,000 iterations.
@@
We conduct experiments on multiple backbone networks. Table VI provides performance metrics for the UPerNet algorithm using different backbones (ViT and LKCA), iteration algorithms, and iteration counts (2000, 4000, 8000, 16000). At each iteration count, LKCA-B consistently outperforms ViT-B in terms of aAcc and mIoU, indicating superior performance of LKCA-B in the UPerNet task. As the iteration count increases, both models show an improvement in performance, but at each iteration count, LKCA-B maintains a leading position. Overall, LKCA-B demonstrates superior performance compared to ViT-B at different iteration counts, with higher pixel accuracy and mean Intersection over Union. Additionally, LKCA-B exhibits advantages in terms of parameter count and computational complexity.
@@
Table VII presents a performance comparison among different backbones (ViT-S, ResNet-50, Swin-S, Twins PCPVT-S, LKCA-S) under various iteration counts and across two different task settings (UPerNet and FPN). In the UPerNet task, LKCA-S consistently outperforms other models at different iteration counts, exhibiting higher aAcc and mIoU. It achieves superior performance compared to other models. Similarly, in the FPN task, LKCA-S demonstrates excellent performance across different iteration counts, showcasing higher aAcc and mIoU, making it competitive against other models. LKCA-S consistently exhibits outstanding performance across different tasks and iteration settings, while maintaining relatively lower computational complexity and parameter counts.
V Conclusion
@@
In this study, we re-examine the relationship between the attention mechanism in visual transformers and the large kernel convolutional networks, proposing a new visual attention called Large Kernel Convolutional Attention (LKCA) aimed for small datasets. It employs a single large kernel convolution to simplify attention operations. In comparison to the previous approach, Large Kernel Attention (LKA), which uses multiple small kernel convolutions, our method achieves the same effects but is more straightforward and simpler. LKCA combines the advantages of convolutional neural networks and visual transformers, featuring a larger receptive field, locality, and parameter sharing. We explained the superiority of LKCA from the perspectives of both convolution and attention, providing equivalent code implementations for each view. Experimental results demonstrate that LKCA, implemented from both convolutional and attention perspectives, exhibits comparable performance. We extensively experimented with LKCA variants of ViT on multiple benchmark datasets, including CIFAR-10, CIFAR-100, SVHN, Tiny-ImageNet, and ADE20K, for classification and segmentation tasks. The experimental outcomes reveal that LKCA demonstrates competitive performance in visual tasks.